# OV-MER: Towards Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ Abstract
Multimodal Emotion Recognition (MER) is a critical research area that seeks to decode human emotions from diverse data modalities. However, existing machine learning methods predominantly rely on predefined emotion taxonomies, which fail to capture the inherent complexity, subtlety, and multi-appraisal nature of human emotional experiences, as demonstrated by studies in psychology and cognitive science. To overcome this limitation, we advocate for introducing the concept of open vocabulary into MER. This paradigm shift aims to enable models to predict emotions beyond a fixed label space, accommodating a flexible set of categories to better reflect the nuanced spectrum of human emotions. To achieve this, we propose a novel paradigm: Open-Vocabulary MER (OV-MER), which enables emotion prediction without being confined to predefined spaces. However, constructing a dataset that encompasses the full range of emotions for OV-MER is practically infeasible; hence, we present a comprehensive solution including a newly curated database, novel evaluation metrics, and a preliminary benchmark. By advancing MER from basic emotions to more nuanced and diverse emotional states, we hope this work can inspire the next generation of MER, enhancing its generalizability and applicability in real-world scenarios.
1 Introduction
Research on emotions has a history spanning two centuries. As early as the 19th century, Charles Darwin conducted pioneering research about the evolutionary origins and possible purposes of emotions, explaining the emotional expressions of humans and animals [1]. In 1884, James revealed the process of emotion generation, noting that stimuli trigger activities in the autonomic nervous system, which in turn produces an emotional experience in the brain [2]. With the rapid development of AI, emotions have garnered increasing attention [3].
The basis of Multimodal Emotion Recognition (MER) lies in the effective modeling of emotions. Current emotion models are primarily categorized into two types: dimensional and discrete models. Dimensional models, particularly those based on psychological theories such as the Circumplex Model of Affect [4], represent emotions within a continuous, multi-dimensional space. The most widely adopted framework uses two or three primary dimensions: valence (the pleasantness-unpleasantness continuum), arousal (the activation-deactivation level), and dominance (the degree of control perceived). These dimensions allow for the quantification of emotional states into measurable numerical values [5]. However, this sophisticated numerical representation demands specialized psychological expertise for accurate interpretation, making it abstract and less descriptive to the general public. This abstraction can result in inconsistencies among different annotators, particularly in complex emotional states that fall between the primary dimensions, thereby complicating subsequent applications and potentially affecting the reliability of emotion recognition systems.
Discrete models, which categorize emotions into distinct classes, tend to mirror the way people naturally perceive and express emotions in daily life. [6] proposed the basic emotion theory, suggesting that there are six basic emotions: anger, disgust, fear, happiness, sadness, and surprise. This theory is widely used in MER, where researchers typically limit the label space to these basic emotions and use multiple annotators to select the most likely label through majority voting. We refer to this task as One-hot MER (OH-MER). Considering that emotions can be compound, researchers further propose Multi-label MER (ML-MER), allowing each sample to have multiple labels [7]. However, both OH-MER and ML-MER generally have limited label spaces. [8] pointed out that humans can express approximately 34,000 distinct emotions. Although some efforts have been made to expand label spaces with more emotional categories, current approaches still fail to capture emotion diversity, inevitably overlooking some of these nuanced emotions.
@@ 1 Introduction
Figure 1: Comparison. (a) Task Comparison: We compare the differences among three tasks (one-hot MER, multi-label MER, and OV-MER) across three aspects (label space, label number, and annotation manner). An in-depth comparison is provided in the Appendix A; (b) Label Comparison: We provide an example to visualize the one-hot and OV labels. More examples are provided in Appendix E. Since the original video contains real people, we use DemoAI to remove personal information to address copyright concerns. In this paper, we use emotion-related descriptions as a bridge to extract OV labels. We observe that OV labels offer a more insightful understanding of the emotional state.
@@
In this paper, we introduce a new MER paradigm, open-vocabulary MER (OV-MER), by introducing the concept of open-vocabulary into MER, enabling the prediction of arbitrary emotion categories. Figure 1(a) provides a comparison between different tasks, and an in-depth comparison is provided in Appendix A. To support this shift, we build a dataset, define evaluation metrics, and develop solutions. (1) Dataset: we propose a human-LLM collaboration strategy to construct the dataset. Compared to human-only annotation, our strategy can leverage LLM to enhance the label richness; (2) Metrics: since there is no fixed label space, the model may predict closely related but differently expressed emotions (e.g., joyful and happy). To provide more reliable evaluation results, we first group similar emotions and specifically design metrics for this task; (3) Solutions: traditional classifiers rely on fixed label spaces. However, OV-MER does not restrict the label space, necessitating the definition of new solutions.
@@ 2 The OV-MERD Dataset Construction
Although the concept of OV-MER is intuitive and holds great promise, its practical implementation faces significant challenges. The main difficulty lies in the broad and subtle range of human emotions, making comprehensive labeling a complex task. Traditional annotation methods are limited by their predefined emotion categories, which are often insufficient for the needs of OV-MER. In Figure 2, we propose a human-LLM collaboration strategy that consists of two steps: CLUE-Multi generation and emotion label extraction. Ultimately, we create a dataset called OV-MERD, which offers a richer set of emotions compared to existing datasets (see Table 1). This dataset is an extension of MER2023 [9]. Specifically, MER2023 is collected from movies and TV series, and we randomly select a portion of data for further annotation to build our OV-MERD.
Table 1: Dataset comparison. See Appendix G for a comprehensive comparison.
@@
MOUD [10] A,V,T Dimensional Emotion 1 1
CMU-MOSI [11] A,V,T Dimensional Emotion 1 1
CH-SIMS [12] A,V,T Dimensional Emotion 1 1
CH-SIMS v2 [13] A,V,T Dimensional Emotion 1 1
SEMAINE [14] A,V,T Dimensional Emotion 5 1
MSP-IMPROV [15] A,V,T Discrete Emotion 4 1
IEMOCAP [16] A,V,T Discrete Emotion 10 1
MELD [17] A,V,T Discrete Emotion 7 1
MER2023 [9] A,V,T Discrete Emotion 6 1
MER2024 [18] A,V,T Discrete Emotion 6 1
@@ 2.1 CLUE-Multi Generation
During the annotation process, we observe that human-LLM collaboration yields more detailed descriptions than the human-only strategy (see Section 5). In this section, we provide a detailed overview of our strategy.
@@
Initially, we attempt to annotate visual and acoustic clues directly. However, the descriptions obtained in this way cannot cover all information. Therefore, we explore using other models for pre-annotation. (1) For video, given the strong visual understanding capabilities of GPT-4V (“gpt-4-vision-preview”), we use it as VLLM for pre-annotation. Since GPT-4V only supports image input, we uniformly sample three frames from each video and input them into GPT-4V. We discuss the reasons for sampling three frames in Appendix J. (2) For audio, we use the open-source SALMONN [19] as ALLM for pre-annotation, as GPT-4V does not support audio input.
@@
As part of our quality assurance procedures, we perform a detailed examination of the pre-annotated results. For visual clues, GPT-4V may generate hallucinated responses, i.e., clues that do not actually exist. Additionally, there are repeated expressions and some temporal association clues are missing. Therefore, we hire annotators to eliminate errors and duplicates, as well as add missing content. For acoustic clues, ALLM struggles to capture emotion-related paralinguistic features. The main reason is that current ALLM mainly focuses on tasks like ASR or audio event detection [19], with less emphasis on paralinguistic information. Hence, we hire multiple annotators to focus on the speaker’s intonation and other emotion-related paralinguistic clues. To reduce subjective bias, we conduct two rounds of manual checks, each involving different annotators. These annotators are experts in affective computing and are familiar with the definitions of emotions. Ultimately, these checked clues can accurately reflect the video content. Appendix K provides the annotation guideline and layout of the annotation platform.
@@
We leverage the reasoning capabilities of LLM to merge all clues. Specifically, we use GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as the LLM and ask it to merge textual, acoustic, and visual clues. The output of this process is an emotion-related description, denoted as CLUE-Multi (see Figure 2). In Appendix J, we further discuss the details of this merging process and the reasons behind it. Overall, the above annotation pipeline reflects the collaboration between humans and LLMs. In Section 5, we compare human-only annotation with our strategy, observing that the latter can generate richer emotions.
@@ 2.2 Ground-truth OV Label Extraction
After that, we use the LLM to extract emotion labels from CLUE-Multi. This process relies on GPT-3.5, which we request to identify emotional states based on the provided descriptions without restricting the label space. See Appendix J for more details.
@@
We further explore the language impact. In Figure 2, we first extract OV labels from English and Chinese descriptions, obtaining $Y_{\text{EE}}$ and $Y_{\text{CC}}$. Then, we translate them into the other language, yielding $Y_{\text{EC}}$ and $Y_{\text{CE}}$. Next, we measure the similarity between different sets and report results in Figure 2. In Appendix M, we detail our metric calculation process. We observe that the labels extracted from different languages exhibit some differences. For example, the similarity score between $Y_{\text{EE}}$ and $Y_{\text{CE}}$ is 0.82, which may be due to the varying definitions of emotions in different languages. To eliminate language influence and achieve consensus labels, we merge the labels extracted from both languages and conduct manual checks. These checked labels are regarded as the ground truth.
@@ 2.3 OV-MERD Dataset
Finally, we construct a dataset called OV-MERD. This dataset is an extension of MER2023 [9], from which we randomly select a portion of samples for further annotation. Table 1 compares OV-MERD with existing datasets. We observe that our OV-MERD dataset contains 236 emotion categories and most samples have 2 to 4 labels, far exceeding those in current datasets. In Appendix H, we observe that OV-MERD encompasses a broader range of emotions, including some that have been rarely discussed in previous research, such as shy, nervous, and grateful.
@@ 3 Evaluation Metric
Table 2: Baseline results. Figure 3 shows the metric calculation process.
@@ 3.1 Grouping
We propose two grouping strategies: one based on GPT and the other based on the emotion wheel (EW) [20]. In the experiments, we use GPT-based grouping by default.
@@
The most direct approach is to use GPT-3.5 to group all labels based on their similarity: Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing emotions with the same meaning. Directly output the results. The output format should be a list containing multiple lists. However, the evaluation results may be affected by the API version. For example, if OpenAI deprecates an old API, the results based on that API will become difficult to reproduce. Additionally, this process is costly (see Appendix N). Therefore, we attempt to find a replacement for GPT-based grouping.
@@
EW is a psychological model that categorizes emotions in a structured manner. The inner part shows core emotions while moving to the outer part reveals more nuanced emotions. Therefore, EW naturally provides emotion grouping information. Since there is no consensus on EW, we select five typical wheels (see Appendix O).
@@ 3.2 Metric Definition
Then, we convert the above emotion grouping information into a function $G(\cdot)$, which can map each label to its group ID. Specifically, suppose $\{y_{i}\}_{i=1}^{M}$ and $\{\hat{y}_{i}\}_{i=1}^{N}$ are the ground truth and predictions, where $M$ and $N$ are the number of labels. We first map each label into its group ID: $\mathcal{Y}=\{G(x)|x\in\{y_{i}\}_{i=1}^{M}\}$ and $\hat{\mathcal{Y}}=\{G(x)|x\in\{\hat{y}_{i}\}_{i=1}^{N}\}$. Then, we design set-based metrics for performance evaluation. Specifically, $\mbox{Precision}_{\mbox{s}}$ indicates the number of correctly predicted labels; $\mbox{Recall}_{\mbox{s}}$ indicates whether the prediction covers all ground truth:
$\mbox{Precision}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\hat{\mathcal{Y}}|},\;\mbox{Recall}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\mathcal{Y}|}.$ (3)
$\mbox{F}_{\mbox{s}}$ is the harmonic mean of two metrics, which is used for the final ranking:
@@
It is important to note that changing the label order in $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ does not result in any score change.

### v3

@@ Abstract
Multimodal Emotion Recognition (MER) is a critical research area that seeks to decode human emotions from diverse data modalities. However, existing machine learning methods predominantly rely on predefined emotion taxonomies, which fail to capture the inherent complexity, subtlety, and multi-appraisal nature of human emotional experiences, as demonstrated by studies in psychology and cognitive science. To overcome this limitation, we advocate for introducing the concept of open vocabulary into MER. This paradigm shift aims to enable models to predict emotions beyond a fixed label space, accommodating a flexible set of categories to better reflect the nuanced spectrum of human emotions. To achieve this, we propose a novel paradigm: Open-Vocabulary MER (OV-MER), which enables emotion prediction without being confined to predefined spaces. However, constructing a dataset that encompasses the full range of emotions for OV-MER is practically infeasible; hence, we present a comprehensive solution including a newly curated database, novel evaluation metrics, and a preliminary benchmark. By advancing MER from basic emotions to more nuanced and diverse emotional states, we hope this work can inspire the next generation of MER, enhancing its generalizability and applicability in real-world scenarios. Code and dataset are available at: https://github.com/zeroQiaoba/AffectGPT.
Machine Learning, ICML
@@
Figure 1: Comparison. (a) Task Comparison: We compare the differences among three tasks (one-hot MER, multi-label MER, and OV-MER) across three aspects (label space, label number, and annotation manner). An in-depth comparison is provided in the Appendix A; (b) Label Comparison: We provide an example to visualize the one-hot and OV labels. More examples are provided in Appendix F. Since the original video contains real people, we use DemoAI to remove personal information to address copyright concerns. In this paper, we use emotion-related descriptions as a bridge to extract OV labels. We observe that OV labels offer a more insightful understanding of the emotional state.
1 Introduction
Research on emotions has a history spanning two centuries. As early as the 19th century, Charles Darwin conducted pioneering research about the evolutionary origins and possible purposes of emotions, explaining the emotional expressions of humans and animals (Darwin, 1872). In 1884, James revealed the process of emotion generation, noting that stimuli trigger activities in the autonomic nervous system, which in turn produces an emotional experience in the brain (James, 1884). With the rapid development of AI, emotions have garnered increasing attention (Minsky, 1988).
The basis of Multimodal Emotion Recognition (MER) lies in the effective modeling of emotions. Current emotion models are primarily categorized into two types: dimensional and discrete models. Dimensional models, particularly those based on psychological theories such as the Circumplex Model of Affect (Russell, 1980), represent emotions within a continuous, multi-dimensional space. The most widely adopted framework uses two or three primary dimensions: valence (the pleasantness-unpleasantness continuum), arousal (the activation-deactivation level), and dominance (the degree of control perceived). These dimensions allow for the quantification of emotional states into measurable numerical values (Warriner et al., 2013). However, this sophisticated numerical representation demands specialized psychological expertise for accurate interpretation, making it abstract and less descriptive to the general public. This abstraction can result in inconsistencies among different annotators, particularly in complex emotional states that fall between the primary dimensions, thereby complicating subsequent applications and potentially affecting the reliability of emotion recognition systems.
Discrete models, which categorize emotions into distinct classes, tend to mirror the way people naturally perceive and express emotions in daily life. Ekman (1992) proposed the basic emotion theory, suggesting that there are six basic emotions: anger, disgust, fear, happiness, sadness, and surprise. This theory is widely used in MER, where researchers typically limit the label space to these basic emotions and use multiple annotators to select the most likely label through majority voting. We refer to this task as One-hot MER (OH-MER). Considering that emotions can be compound, researchers further propose Multi-label MER (ML-MER), allowing each sample to have multiple labels (Li et al., 2017). However, both OH-MER and ML-MER generally have limited label spaces. Plutchik (2001) pointed out that humans can express approximately 34,000 distinct emotions. Although some efforts have been made to expand label spaces with more emotional categories, current approaches still fail to capture emotional diversity, inevitably overlooking some of these nuanced emotions.
@@ 1 Introduction
In this paper, we introduce a new MER paradigm, open-vocabulary MER (OV-MER), by introducing the concept of open-vocabulary into MER, enabling the prediction of arbitrary emotion categories. Figure 1(a) provides a comparison between different tasks, and an in-depth comparison is provided in Appendix A. To support this shift, we build a dataset, define evaluation metrics, and develop solutions. (1) Dataset: we propose a human-LLM collaboration strategy to construct the dataset. Compared to human-only annotation, our strategy can leverage LLM to enhance the label richness; (2) Metrics: since there is no fixed label space, the model may predict closely related but differently expressed emotions (e.g., joyful and happy). To provide more reliable evaluation results, we first group similar emotions and specifically design metrics for this task; (3) Solutions: traditional discriminative classifiers rely on fixed label spaces. However, OV-MER does not restrict the label space, necessitating the definition of new solutions.
@@ 2 The OV-MERD Dataset Construction
Although the concept of OV-MER is intuitive and holds great promise, its practical implementation faces significant challenges. The main difficulty lies in the broad and subtle range of human emotions, making comprehensive labeling a complex task. Traditional annotation methods are limited by their predefined emotion categories, which are often insufficient for the needs of OV-MER. In Figure 2, we propose a human-LLM collaboration strategy that consists of two steps: CLUE-Multi generation and emotion label extraction. Ultimately, we create a dataset, OV-MERD, which offers a richer set of emotions compared to existing datasets (see Table 1). This dataset is an extension of MER2023 (Lian et al., 2023). Specifically, MER2023 is collected from movies and TV series, with most samples consisting of single-person videos featuring relatively complete speech content. The use of this dataset has been approved by the dataset owners. We randomly selected a subset of MER2023 for further annotation to construct our OV-MERD dataset. Additional details about MER2023 can be found in Appendix E.
Table 1: Dataset comparison. See Appendix H for a comprehensive comparison.
@@
MOUD (Pérez-Rosas et al., 2013) A,V,T Dimensional Emotion 1 1
CMU-MOSI (Zadeh et al., 2017) A,V,T Dimensional Emotion 1 1
CH-SIMS (Yu et al., 2020) A,V,T Dimensional Emotion 1 1
CH-SIMS v2 (Liu et al., 2022b) A,V,T Dimensional Emotion 1 1
SEMAINE (McKeown et al., 2011) A,V,T Dimensional Emotion 5 1
MSP-IMPROV (Busso et al., 2016) A,V,T Discrete Emotion 4 1
IEMOCAP (Busso et al., 2008) A,V,T Discrete Emotion 10 1
MELD (Poria et al., 2019) A,V,T Discrete Emotion 7 1
MER2023 (Lian et al., 2023) A,V,T Discrete Emotion 6 1
MER2024 (Lian et al., 2024a) A,V,T Discrete Emotion 6 1
@@ 2.1 CLUE-Multi Generation
During the annotation process, we observe that human-LLM collaboration yields more detailed descriptions than the human-only strategy (see Section 5). In this section, we provide a detailed overview of our strategy. For manual checks, to maintain high-quality annotations, all annotators must pass a preliminary test. This test evaluates their performance on 12 samples, each of which was previously annotated by five annotators with full agreement. Annotators who perform poorly are removed from the annotator pool. More annotation details can be found in Appendix L.
@@
Initially, we attempt to annotate visual and acoustic clues directly. However, the descriptions obtained in this way cannot cover all information. Therefore, we explore using other models for pre-annotation. (1) For video, given the strong visual understanding capabilities of GPT-4V (“gpt-4-vision-preview”), we use it as VLLM for pre-annotation. Since GPT-4V only supports image input, we uniformly sample three frames from each video and input them into GPT-4V. We discuss the reasons for sampling three frames in Appendix K. (2) For audio, we use the open-source SALMONN (Tang et al., 2023) as ALLM for pre-annotation, as GPT-4V does not support audio input.
@@
As part of our quality assurance procedures, we perform a detailed examination of the pre-annotated results. For visual clues, GPT-4V may generate hallucinated responses, i.e., clues that do not actually exist. Additionally, there are repeated expressions and some temporal association clues are missing. Therefore, we hire annotators to eliminate errors and duplicates, as well as add missing content. For acoustic clues, ALLM struggles to capture emotion-related paralinguistic features. The main reason is that current ALLM mainly focuses on tasks like ASR or audio event detection (Tang et al., 2023), with less emphasis on paralinguistic information. Hence, we hire multiple annotators to focus on the speaker’s intonation and other emotion-related paralinguistic clues. To reduce subjective bias, we conduct two rounds of manual checks. Ultimately, these checked clues can accurately reflect the video content. Appendix L provides the annotation guideline and layout of the annotation platform.
@@
We leverage the reasoning capabilities of LLM to merge all clues. Specifically, we use GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as the LLM and ask it to merge textual, acoustic, and visual clues. The output is an emotion-related description, denoted as CLUE-Multi (see Figure 2). It is worth noting that we did not perform additional manual checks of the generated CLUE-Multi, as GPT-3.5 consistently produced reasonable and logical results. This reliability likely stems from the GPT-series models’ exceptional performance in reading comprehension (Brown et al., 2020) (close to human performance), where multi-clue integration is a core functionality. Therefore, we skip the manual inspection of CLUE-Multi, striking a balance between dataset reliability and construction efficiency. In Appendix K, we discuss the details of this merging process and the reasons behind it. The above annotation pipeline reflects the collaboration between humans and LLMs.
@@ 2.2 Ground-truth OV Label Extraction
After that, we use the LLM to extract emotion labels from CLUE-Multi. This process relies on GPT-3.5, which we request to identify emotional states based on the provided descriptions without restricting the label space. See Appendix K for more details.
@@
We further explore the language impact. In Figure 2, we first extract OV labels from English and Chinese descriptions, obtaining $Y_{\text{EE}}$ and $Y_{\text{CC}}$. Then, we translate them into the other language, yielding $Y_{\text{EC}}$ and $Y_{\text{CE}}$. Next, we measure the similarity between different sets and report results in Figure 2. In Appendix N, we detail our metric calculation process. We observe that the labels extracted from different languages exhibit some differences. For example, the similarity score between $Y_{\text{EE}}$ and $Y_{\text{CE}}$ is 0.82, which may be due to the varying definitions of emotions in different languages. To eliminate language influence and achieve consensus labels, we merge the labels extracted from both languages and conduct manual checks. These checked labels are regarded as the ground truth.
@@ 2.3 OV-MERD Dataset
Finally, we construct a dataset called OV-MERD. This dataset is an extension of MER2023 (Lian et al., 2023), from which we randomly select a portion of samples for further annotation. Table 1 compares OV-MERD with existing datasets. We observe that our OV-MERD dataset contains 236 emotion categories, and most samples have 2 to 4 labels, far exceeding those in current datasets. In Appendix I, we observe that OV-MERD encompasses a broader range of emotions, including some that have been rarely discussed in previous research, such as shy, nervous, and grateful.
@@ 3 Evaluation Metric
Table 2: Main results. Figure 3 illustrates the metric calculation process. The primary distinction between CLUE-MLLM (Baselines) and CLUE-M/A/T/V (Upper-Bound Performance) is whether manually verified clues are utilized.
@@ 3.1 Grouping
We propose two grouping strategies: one based on GPT and the other based on the emotion wheel (EW) (Plutchik, 1980). In the experiments, we use GPT-based grouping by default.
@@
The most direct approach is to use GPT-3.5 to group all labels based on their similarity: Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing emotions with the same meaning. Directly output the results. The output format should be a list containing multiple lists. However, the evaluation results may be affected by the API version. For example, if OpenAI deprecates an old API, the results based on that API will become difficult to reproduce. Additionally, this process is costly (see Appendix O). Therefore, we attempt to find a replacement for GPT-based grouping.
@@
EW is a psychological model that categorizes emotions in a structured manner. The inner part shows core emotions, while moving to the outer part reveals more nuanced emotions. Therefore, EW naturally provides emotion grouping information. Since there is no consensus on EW, we select five typical wheels (see Appendix P).
@@ 3.2 Metric Definition
Then, we convert the above emotion grouping information into a function $G(\cdot)$, which can map each label to its group ID. Specifically, suppose $\{y_{i}\}_{i=1}^{M}$ and $\{\hat{y}_{i}\}_{i=1}^{N}$ are the ground truth and predictions, where $M$ and $N$ are the number of labels. We first map each label into its group ID: $\mathcal{Y}=\{G(x)|x\in\{y_{i}\}_{i=1}^{M}\}$ and $\hat{\mathcal{Y}}=\{G(x)|x\in\{\hat{y}_{i}\}_{i=1}^{N}\}$. Then, we design set-based metrics for performance evaluation. Specifically, $\mbox{Precision}_{\mbox{s}}$ indicates the number of correctly predicted labels; $\mbox{Recall}_{\mbox{s}}$ indicates whether the prediction covers all ground truth; $\mbox{F}_{\mbox{s}}$ is the harmonic mean of two metrics, which is used for the final ranking:
$\mbox{Precision}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\hat{\mathcal{Y}}|},\;\mbox{Recall}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\mathcal{Y}|},$ (3)
@@
It is important to note that changing the label order in $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ does not result in any score change. The subscript “s” indicates that these metrics are set-based, distinguishing them from the traditional single-label metrics.